OpenStudy (anonymous):

−3a8=6 x−2+16=11 y−94+6=3 5b−8=7b+12 3(2a+4)=2(a+3)

OpenStudy (benlindquist):

add you like terms, or multiply your like terms

OpenStudy (anonymous):

Linear Equation Assessment 9-12= 3, 6-9=2, 0-6=1